# Distal Causal Excursion Effects: Modeling Long-Term Effects of Time-Varying Treatments in Micro-Randomized Trials

### v1

@@ 2.3 Causal Assumptions and Identification
Positivity and sequential ignorability are guaranteed by the MRT design. SUTVA holds for most MRTs without the social aspect but will be violated if interference is present, that is, if the treatment assigned to one participant affects the potential outcome of another participant. In such settings, a framework that incorporates causal interference is needed (Hudgens & Halloran, 2008; Shi et al., 2022). We do not consider such settings here.
Let $p^{1}_{t}(a\mid H_{t})$ denote the probability mass function of $A_{t}$ conditional on $H_{t}$ under the decision rule $d_{t}^{1}$: $p^{1}_{t}(a\mid H_{t})=P_{A_{t}\sim d_{t}^{1}(H_{t})}(A_{t}=a\mid H_{t})$, i.e., $p^{1}_{t}(1\mid H_{t})=I_{t}$ and $p^{1}_{t}(0\mid H_{t})=1-I_{t}$. Similarly, let $p^{0}_{t}(a\mid H_{t})$ denote the distribution of $A_{t}$ under $d_{t}^{0}$, i.e., $p^{0}_{t}(1\mid H_{t})=0$ and $p^{0}_{t}(0\mid H_{t})=1$. In the Supplementary Material B, we prove the following identification result for $\tau(t,s)$.
Theorem 1.
@@ 3 Robust Estimators for DCEE
Let $f(t,s)$ be a pre-specified $p$-dimensional feature vector that depends on the decision point index and the effect modifier value. We consider estimating the best linear projection of $\tau(t,s)$ on $f(t,s)$, averaged over all decision points. Specifically, the true parameter $\beta^{\star}\in\mathbb{R}^{p}$ is defined as
@@
where $\omega(t)$ is a pre-specified weight function satisfying $\sum_{t=1}^{T}\omega(t)=1$. This is similar to linearly parameterizing $\tau(t,S_{t})=f(t,S_{t})^{T}\beta^{\star}$, but by defining $\beta^{\star}$ this way the estimator is interpretable even when the linear parameterization does not reflect the true data-generating distribution.
Researchers may set different $f(t,s)$ and different $\omega(t)$ based on the specific scientific question. For example, when $S_{t}=\emptyset$, setting $f(t,s)=(1,t)$ or $f(t,s)=(1,t,t^{2})$ models how the DCEE may vary linearly or quadratically over the decision point where the excursion takes place. One may also include basis functions of $t$ in $f(t,s)$ to allow flexibility (see Section 5 for such an example). When $S_{t}\neq\emptyset$, $f(t,s)$ can be either additive in $t,s$ or include their interaction. Regarding the weights $\omega(t)$, setting $\omega(t)=1/T$ for all $t$ ensures that all decision points contribute to the estimation of $\beta$ equally. However, if the goal is to estimate $\tau(t,s)$ at a particular $t=t_{0}$, one can set $\omega(t_{0})=1$ and $\omega(t)=0$ for all $t\neq t_{0}$.
@@
$\beta^{\star}=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}(I_{t})\mathbb{E}\Big{\{}\mathbb{E}(Y|H_{t},A_{t}=1,I_{t}=1)-\mathbb{E}(Y|H_{t},A_{t}=1,I_{t}=0)\Big{\}}.$ (7)
@@
×$\times\sum_{t=1}^{T}\mathbb{E}(I_{t})\mathbb{E}\Big{[}\big{\{}\mathbb{E}(Y|H_{t},A_{t}=1,I_{t}=1)-\mathbb{E}(Y|H_{t},A_{t}=1,I_{t}=0)\big{\}}f(t,S_{t})\Big{]}.$ (8)
@@
The identification equation (3) motivates the following rudimentary estimating function for $\beta$:
$\xi(\beta):=\sum_{t=1}^{T}\omega(t)\bigg{\{}\frac{p_{t}^{1}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}Y-\frac{p_{t}^{0}(A_{t}\mid H_{t})}{p_{t}(A_{t}\mid H_{t})}Y-f(t,S_{t})^{T}\beta\bigg{\}}f(t,S_{t}).$ (9)
We further subtract from $\xi(\beta)$ its projection on the score functions of the treatment selection probabilities to obtain a more efficient estimating function (Robins, 1999):
@@
The form of $\phi(\beta,\mu)$ motivates two estimators, $\hat{\beta}$ and $\tilde{\beta}$, as depicted in Algorithms 1 and 2. $\hat{\beta}$ does not use cross-fitting and $\tilde{\beta}$ uses cross-fitting (Chernozhukov et al., 2018).
Stage 1: Fit $\mathbb{E}(Y\mid H_{t},A_{t})$ for $t\in[T]$. Denote the fitted model by $\hat{\mu}_{t}(H_{t},A_{t})$. Denote $\hat{\mu}:=(\hat{\mu}_{1},\ldots,\hat{\mu}_{T})$.
@@
Stage 2: Obtain $\hat{\beta}$ by solving $\mathbb{P}_{n}\phi(\beta,\hat{\mu})=0$ with $\phi$ defined in (10).
@@
Algorithm 1 A two-stage estimator $\hat{\beta}$ (without cross-fitting)
@@
Stage 1: Take a $K$-fold equally-sized random partition $(B_{k})_{k=1}^{K}$ of observation indices $[n]=\{1,\ldots,n\}$. Define $B_{k}^{c}=[n]\setminus B_{k}$ for $k\in[K]$. For each $k\in[K]$, use solely observations from $B_{k}^{c}$ and fit $\mathbb{E}(Y\mid H_{t},A_{t})$ for $t\in[T]$. The fitted models using $B_{k}^{c}$ are denoted by $\hat{\mu}_{kt}$, and let $\hat{\mu}_{k}:=(\hat{\mu}_{k1},\ldots,\hat{\mu}_{kT})$.
@@
Stage 2: Obtain $\tilde{\beta}$ by solving $K^{-1}\sum_{k=1}^{K}\mathbb{P}_{n,k}\phi(\beta,\hat{\mu}_{k})=0$. Here $\mathbb{P}_{n,k}$ denotes empirical average over observations from $B_{k}$.
@@
Algorithm 2 A two-stage estimator $\tilde{\beta}$ (with cross-fitting)
@@
We establish the asymptotic normality for $\hat{\beta}$ and $\tilde{\beta}$ in Theorem 2, which is proved in the Supplementary Material E.
Theorem 2 (Asymptotic normality of $\hat{\beta}$ and $\tilde{\beta}$).
@@
Suppose Assumption 1 hold and consider $\beta^{\star}$ defined in (6). Suppose $\hat{\mu}$ converges to some limit $\mu^{\prime}$ (not necessarily the true $\mu^{\star}$) in $L_{2}$. Under regularity conditions, we have
$\sqrt{n}(\hat{\beta}-\beta^{\star})\stackrel{{\scriptstyle d}}{{\to}}N(0,V)\text{ as }n\to\infty,$ (11)
@@
with $V:=\mathbb{E}\{\partial_{\beta}\phi(\beta^{\star},\mu^{\prime})\}^{-1}\mathbb{E}[\{\phi(\beta^{\star},\mu^{\prime})\phi(\beta^{\star},\mu^{\prime})^{T}\}]\mathbb{E}\{\partial_{\beta}\phi(\beta^{\star},\mu^{\prime})\}^{-1,T}$. $V$ can be consistently estimated by $\mathbb{P}_{n}\{\partial_{\beta}\phi(\hat{\beta},\hat{\mu})\}^{-1}\mathbb{P}_{n}[\{\phi(\hat{\beta},\hat{\mu})\phi(\hat{\beta},\hat{\mu})^{T}\}]\mathbb{P}_{n}\{\partial_{\beta}\phi(\hat{\beta},\hat{\mu})\}^{-1,T}$.
@@
For the estimator $\tilde{\beta}$ that uses cross-fitting, assume that $\hat{\mu}_{k}$ converges to some limit $\mu^{\prime}$ (not necessarily the true $\mu^{\star}$) in $L_{2}$ for each $k\in[K]$. Under regularity conditions, the asymptotic normality (11) holds with $\hat{\beta}$ replaced by $\tilde{\beta}$, in which case $V$ can be consistently estimated by
$\bigg{[}\frac{1}{K}\sum_{k=1}^{K}\mathbb{P}_{n,k}\big{\{}\partial_{\beta}\phi(\hat{\beta},\hat{\mu}_{k})\big{\}}\bigg{]}^{-1}\bigg{[}\frac{1}{K}\sum_{k=1}^{K}\mathbb{P}_{n,k}\big{\{}\phi(\hat{\beta},\hat{\mu}_{k})\phi(\hat{\beta},\hat{\mu}_{k})^{T}\big{\}}\bigg{]}\bigg{[}\frac{1}{K}\sum_{k=1}^{K}\mathbb{P}_{n,k}\big{\{}\partial_{\beta}\phi(\hat{\beta},\hat{\mu}_{k})\big{\}}\bigg{]}^{-1,T},$
@@
$\hat{\beta}$ and $\tilde{\beta}$ are robust in the sense that their consistency and asymptotical normality hold with arbitrarily fitted nuisance parameter $\hat{\mu}$ that may not need to converge to the true $\mu^{\star}$. Furthermore, the asymptotic variance $V$ only involves $\mu^{\prime}$, the limit of $\hat{\mu}$, and otherwise does not depend on how $\hat{\mu}$ is fitted. This is because the estimating function $\phi(\beta,\mu)$ is globally robust (Cheng et al., 2023), which we prove in the Supplementary Material E. Of course, a better $\hat{\mu}$ will enhance the efficiency of $\hat{\beta}$ and $\tilde{\beta}$. Using nuisance parameters that are not required for identification to improve efficiency is a common technique in causal inference (e.g., Tsiatis et al., 2008; Lok, 2024; Van Lancker et al., 2024). ∎
@@ 4.1 Generative Model and True Parameter Value
We construct a generative model that simultaneously incorporates the following common complications in actual MRTs: endogenous time-varying covariates that are influenced by past treatments and affect future treatment assignments, effect moderation by time-varying covariates, current treatments interacting with past treatments which reflects user burden, and eligibility constraints.
In the generative model, we set the total number of decision points per person to $T=30$. Data from different individuals are i.i.d. We describe below the data-generating process for a generic individual. $X_{t}$ is a time-varying continuous covariate that depends on $A_{t-1}$ and $X_{t-1}$: $X_{t}=\theta_{0}+\theta_{1}A_{t-1}+\theta_{2}X_{t-1}+\eta_{t}$ with $\eta_{t}\sim N(0,1)$ being an exogenous error. $Z_{t}$ is a time-varying binary covariate that depends on $A_{t-1}$ and $Z_{t-1}$: $Z_{t}\sim\text{Bernoulli}\{\text{expit}(\zeta_{0}+\zeta_{1}A_{t-1}+\zeta_{2}Z_{t-1})\}$. We set $X_{0}=Z_{0}=A_{0}=0$. The eligibility indicator is generated exogenously: $I_{t}\sim\text{Bernoulli}(0.8)$. The randomization probability for $A_{t}$ is $P(A_{t}=1\mid H_{t},I_{t}=1)=\text{expit}\{(t-T/2)/T+Z_{t}-0.5+X_{t}/6\}$, where the transformations of the variables make their influence on $A_{t}$ comparable. The outcome is generated by
@@
where $g(\cdot)$ is nonlinear and equals the probability density function for $\text{Beta}(2,2)$ distribution. The observed data for a generic individual is $\{X_{t},Z_{t},I_{t},A_{t}:1\leq t\leq T\}\cup\{Y\}$, and the randomization probability $P(A_{t}=1\mid H_{t},I_{t}=1)$ is known as part of the trial design.
The parameter values are set as follows: $\theta_{0}=-0.5$, $\theta_{1}=\theta_{2}=0.5$; $\zeta_{0}=-1$, $\zeta_{1}=\zeta_{2}=1$; $\alpha_{t}=1+2(t-1)/(T-1)$; $\nu_{t}=1+(t-1)/(T-1)$; $\gamma_{t}=1+0.5(t-1)/(T-1)$; $\lambda_{t}=-1-(t-1)/(T-1)$; $\xi_{t}=1+(t-1)/(T-1)$. The $\lambda_{t}$ takes negative values, reflecting user burden (Example 2). The magnitudes of $\alpha_{t},\nu_{t},\gamma_{t},\lambda_{t},\xi_{t}$ all increase with $t$, characterizing a scenario where the variables more proximal to the distal outcome $Y$ have a larger impact on $Y$.
We consider two sets of estimands. We set $\omega(t)=1/T$. The first estimand corresponds to the fully marginal DCEE by setting $S_{t}=\emptyset$ in (2) and $f(t,S_{t})=1$ in (6), and the true parameter value is $\beta^{\star}_{1}=1.603$. The second set of estimands corresponds to the DCEE moderated by the binary covariate $Z_{t}$, i.e., through setting $S_{t}=Z_{t}$ and $f(t,S_{t})=(1,Z_{t})^{T}$. The true parameter values are $(\beta^{\star}_{2},\beta^{\star}_{3})=(1.207,0.881)$. The numerical computation of $\beta^{\star}_{1}$ and $(\beta^{\star}_{2},\beta^{\star}_{3})$ is detailed in the Supplementary Material F.

### v2

@@ 2.3 Causal Assumptions and Identification
Positivity and sequential ignorability are guaranteed by the MRT design. SUTVA holds for most MRTs without the social aspect but is violated under interference, where one participant’s treatment affects another’s outcome. In such cases, frameworks incorporating causal interference are needed (Hudgens & Halloran, 2008; Shi et al., 2022), which we do not consider here. In Supplementary Material B, we prove the following identification result for $\tau(t,s)$.
Theorem 2.1.
@@
$\tau(t,s)$ $=\mathbb{E}\bigg{\{}\frac{{\mathds{1}}(A_{t}=I_{t})}{P(A_{t}=I_{t}\mid H_{t})}Y-\frac{{\mathds{1}}(A_{t}=0)}{P(A_{t}=0\mid H_{t})}Y~\bigg{|}~S_{t}=s\bigg{\}}.$ (3)
$=\mathbb{E}\Big{\{}\mathbb{E}(Y\mid H_{t},A_{t}=I_{t})-\mathbb{E}(Y\mid H_{t},A_{t}=0)\mid S_{t}=s\Big{\}}$ (4)
@@ 3 Robust Estimators for DCEE
Let $f(t,s)$ be a pre-specified $p$-dimensional feature vector that depends on the decision point index and the effect modifier value. We consider estimating the best linear projection of $\tau(t,s)$ on $f(t,s)$, averaged over decision points. Specifically, the true parameter $\beta^{\star}\in\mathbb{R}^{p}$ is defined as
@@
where $\omega(t)$ is a pre-specified weight function with $\sum_{t=1}^{T}\omega(t)=1$. This is similar to linearly parameterizing $\tau(t,S_{t})=f(t,S_{t})^{T}\beta^{\star}$, but by defining $\beta^{\star}$ this way ensures interpretability even when the linear model is misspecified.
Researchers can choose different $f(t,s)$ and $\omega(t)$ based on the scientific question. For example, when $S_{t}=\emptyset$, setting $f(t,s)=(1,t)$ or $(1,t,t^{2})$ models how the DCEE varies linearly or quadratically over the decision point where the excursion takes place. Basis functions of $t$ can also be included for flexibility (see Section 5). When $S_{t}\neq\emptyset$, $f(t,s)$ can be additive in $t,s$ or include interactions. For weights $\omega(t)$, setting $\omega(t)=1/T$ equally weights all decision points in estimating $\beta$. On the other hand, if the goal is to estimate $\tau(t,s)$ at a particular $t=t_{0}$, one can set $\omega(t_{0})=1$ and $\omega(t)=0$ for all $t\neq t_{0}$.
@@
$\beta^{\star}=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}(I_{t})\mathbb{E}\Big{\{}\mathbb{E}(Y|H_{t},I_{t}=1,A_{t}=1)-\mathbb{E}(Y|H_{t},I_{t}=1,A_{t}=0)\Big{\}}.$ (7)
@@
$\times\sum_{t=1}^{T}\mathbb{E}(I_{t})\mathbb{E}\Big{[}\big{\{}\mathbb{E}(Y|H_{t},A_{t}=1,I_{t}=1)-\mathbb{E}(Y|H_{t},A_{t}=1,I_{t}=0)\big{\}}f(t,S_{t})\Big{]}.$ (8)
@@
The identification equation (3) motivates a preliminary estimating function for $\beta$:
$\xi(\beta):=\sum_{t=1}^{T}\omega(t)\bigg{\{}\frac{{\mathds{1}}(A_{t}=I_{t})}{P(A_{t}=I_{t}\mid H_{t})}Y-\frac{{\mathds{1}}(A_{t}=0)}{P(A_{t}=0\mid H_{t})}Y-f(t,S_{t})^{T}\beta\bigg{\}}f(t,S_{t}).$ (9)
We further subtract from $\xi(\beta)$ its projection on the score functions of the treatment selection probabilities to obtain a more efficient estimating function (Robins, 1999b):
@@
The form of $\phi(\beta,\mu)$ motivates two estimators, $\widehat{\beta}$ and $\widetilde{\beta}$, as depicted in Algorithms 1 and 2. $\widehat{\beta}$ does not use cross-fitting and $\widetilde{\beta}$ uses cross-fitting (Chernozhukov et al., 2018).
Stage 1: Fit $\mathbb{E}(Y\mid H_{t},A_{t})$ for $t\in[T]$. Denote the fitted model by $\widehat{\mu}_{t}(H_{t},A_{t})$. Denote $\widehat{\mu}:=(\widehat{\mu}_{1},\ldots,\widehat{\mu}_{T})$.
@@
Stage 2: Obtain $\widehat{\beta}$ by solving $\mathbb{P}_{n}\phi(\beta,\widehat{\mu})=0$ with $\phi$ defined in (10).
@@
Algorithm 1 A two-stage estimator $\widehat{\beta}$ (without cross-fitting)
@@
Stage 1: Take a $K$-fold equally-sized random partition $(B_{k})_{k=1}^{K}$ of observation indices $n]=\{1,\ldots,n\}$. Define $B_{k}^{c}=[n]\setminus B_{k}$ for $k\in[K]$. For each $k\in[K]$, use solely observations from $B_{k}^{c}$ and fit $\mathbb{E}(Y\mid H_{t},A_{t})$ for $t\in[T]$. The fitted models using $B_{k}^{c}$ are denoted by $\widehat{\mu}_{kt}$, and let $\widehat{\mu}_{k}:=(\widehat{\mu}_{k1},\ldots,\widehat{\mu}_{kT})$.
@@
Stage 2: Obtain $\widetilde{\beta}$ by solving $K^{-1}\sum_{k=1}^{K}\mathbb{P}_{n,k}\phi(\beta,\widehat{\mu}_{k})=0$. Here $\mathbb{P}_{n,k}$ denotes empirical average over observations from $B_{k}$.
@@
Algorithm 2 A two-stage estimator $\widetilde{\beta}$ (with cross-fitting)
@@
We establish the asymptotic normality for $\widehat{\beta}$ and $\widetilde{\beta}$ in Theorem 3.1, which is proved in the Supplementary Material E.
Theorem 3.1 (Asymptotic normality of $\widehat{\beta}$ and $\widetilde{\beta}$).
@@
Suppose Assumption 1 hold and consider $\beta^{\star}$ defined in (6). Suppose $\widehat{\mu}$ converges to some limit $\mu^{\prime}$ (not necessarily the true $\mu^{\star}$) in $L_{2}$. Under regularity conditions, we have
[…]
@@
with $\mathbb{E}\{\partial_{\beta}\phi(\beta^{\star},\mu^{\prime})\}^{-1}\mathbb{E}[\{\phi(\beta^{\star},\mu^{\prime})\phi(\beta^{\star},\mu^{\prime})^{T}\}]\mathbb{E}\{\partial_{\beta}\phi(\beta^{\star},\mu^{\prime})\}^{-1,T}. VV$ can be consistently estimated by $\mathbb{P}_{n}\{\partial_{\beta}\phi(\widehat{\beta},\widehat{\mu})\}^{-1}\mathbb{P}_{n}[\{\phi(\widehat{\beta},\widehat{\mu})\phi(\widehat{\beta},\widehat{\mu})^{T}\}]\mathbb{P}_{n}\{\partial_{\beta}\phi(\widehat{\beta},\widehat{\mu})\}^{-1,T}$.
@@
For the estimator $\widetilde{\beta}$ that uses cross-fitting, assume that $\widehat{\mu}_{k}$ converges to some limit $\mu^{\prime}$ (not necessarily the true $\mu^{\star}$) in $L_{2}$ for each $k\in[K]$. Under regularity conditions, the asymptotic normality (11) holds with $\widehat{\beta}$ replaced by $\widetilde{\beta}$, in which case $V$ can be consistently estimated by
$\bigg{[}\frac{1}{K}\sum_{k=1}^{K}\mathbb{P}_{n,k}\big{\{}\partial_{\beta}\phi(\widehat{\beta},\widehat{\mu}_{k})\big{\}}\bigg{]}^{-1}\bigg{[}\frac{1}{K}\sum_{k=1}^{K}\mathbb{P}_{n,k}\big{\{}\phi(\widehat{\beta},\widehat{\mu}_{k})\phi(\widehat{\beta},\widehat{\mu}_{k})^{T}\big{\}}\bigg{]}\bigg{[}\frac{1}{K}\sum_{k=1}^{K}\mathbb{P}_{n,k}\big{\{}\partial_{\beta}\phi(\widehat{\beta},\widehat{\mu}_{k})\big{\}}\bigg{]}^{-1,T},$
@@
$\widehat{\beta}$ and $\widetilde{\beta}$ are robust in the sense that their consistency and asymptotical normality hold even with arbitrarily fitted nuisance parameter $\widehat{\mu}$, which need not converge to the true $\mu^{\star}$. Furthermore, the asymptotic variance $V$ involves only $\mu^{\prime}$, the limit of $\widehat{\mu}$, and does not depend on how $\widehat{\mu}$ is fitted. This is because the estimating function $\phi(\beta,\mu)$ is globally robust (Cheng et al., 2023), as we prove in Supplementary Material E. Of course, a better $\widehat{\mu}$ improves the efficiency of $\widehat{\beta}$ and $\widetilde{\beta}$. Using nuisance parameters not required for identification to improve efficiency is common in causal inference (e.g., Tsiatis et al., 2008; Lok, 2024). ∎
@@ 4.1 Generative Model and True Parameter Value
We construct a generative model that incorporates common complications in actual MRTs: endogenous time-varying covariates influenced by past treatments and affecting future treatments, effect moderation by time-varying covariates, interactions between current and past treatments that reflect user burden, and eligibility constraints.
In the generative model, each individual has $T=30$ decision points, and data across individuals are i.i.d. For a generic individual, $X_{t}$ is a time-varying continuous covariate that depends on $A_{t-1}$ and $X_{t-1}$: $X_{t}=\theta_{0}+\theta_{1}A_{t-1}+\theta_{2}X_{t-1}+\eta_{t}$, with $\eta_{t}\sim N(0,1)$ exogenous. $Z_{t}$ is a binary time-varying covariate that depends on $A_{t-1}$ and $Z_{t-1}$: $Z_{t}\sim\text{Bernoulli}\{\text{expit}(\zeta_{0}+\zeta_{1}A_{t-1}+\zeta_{2}Z_{t-1})\}$. We set $X_{0}=Z_{0}=A_{0}=0$. The eligibility indicator is exogenous: $I_{t}\sim\text{Bernoulli}(0.8)$. The randomization probability is $P(A_{t}=1\mid H_{t},I_{t}=1)=\text{expit}\{(t-T/2)/T+Z_{t}-0.5+X_{t}/6\}$, ensuring comparable influence of each variable on $A_{t}$. The outcome is
@@
where $g(\cdot)$ is the $\text{Beta}(2,2)$ density function (nonlinear), and $\epsilon\sim N(0,1)$ is exogeneous. The observed data for an individual is $\{X_{t},Z_{t},I_{t},A_{t}:1\leq t\leq T\}\cup\{Y\}$, with $P(A_{t}=1\mid H_{t},I_{t}=1)$ is known from the trial design.
The parameter values are set as follows: $\theta_{0}=-0.5$, $\theta_{1}=\theta_{2}=0.5$; $\zeta_{0}=-1$, $\zeta_{1}=\zeta_{2}=1$; $\alpha_{t}=1+2(t-1)/(T-1)$; $\nu_{t}=1+(t-1)/(T-1)$; $\gamma_{t}=1+0.5(t-1)/(T-1)$; $\lambda_{t}=-1-(t-1)/(T-1)$; $\xi_{t}=1+(t-1)/(T-1)$. The negative $\lambda_{t}$ reflects user burden (Example 2). The magnitudes of $\alpha_{t},\nu_{t},\gamma_{t},\lambda_{t},\xi_{t}$ all increase with $t$, characterizing a scenario where variables closer to the distal outcome $Y$ have a larger impact on $Y$.
We consider two sets of estimands and set $\omega(t)=1/T$. The first estimand corresponds to the fully marginal DCEE by setting $S_{t}=\emptyset$ in (2) and $f(t,S_{t})=1$ in (6), and the true parameter value is $\beta^{\star}_{1}=1.603$. The second set of estimands corresponds to the DCEE moderated by the binary covariate $Z_{t}$, i.e., through setting $S_{t}=Z_{t}$ and $f(t,S_{t})=(1,Z_{t})^{T}$. The true parameter values are $(\beta^{\star}_{2},\beta^{\star}_{3})=(1.207,0.881)$. The numerical computation of $\beta^{\star}_{1}$ and $(\beta^{\star}_{2},\beta^{\star}_{3})$ is detailed in Supplementary Material F.